Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Joint Operation and Readiness Logbook, Contact: T.Haas, A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 575
Entry time: 05 Nov 2018, 12:25
|
Author: |
Adriano Violante |
Location: |
General |
Severity: |
Information |
Subject: |
Status report after dispatch meeting 09.30 |
|
|
SPB
FXE
SQS
- Crates for Loop12 have been installed. Loop will probably run by today or tomorrow.
- Loop 2 and 3 need to be updated because of interlock definitions.
- Issue with valves behavior in Loop 3 and 5. AE will check.
- Encoder w/o 5 V. BB checks.
- Pfeiffer Motor valve issue will be fixed by next week.
SCS
MID
- AGIPD will move in the hutch this week. Preparation ongoing.
- Loop11 will be ready early next week
HED
XPD
- Several tests planned for the end of the week, including the feedback device by CAS. PRCs will be made aware of the activities in the tunnels.
VAC
AE
- PPL update in SASE1 today.
- Update of the Loop 5 (AGIPD) tomorrow
ITDM
CAS
- Issues in the calibration pipeline will be coordinated with the relevant groups (DET, AE) before and during the next chapter's meeting.
AoB
|